Circular No. 8771

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2006lu–2006mx Several apparent supernovae have been found on unfiltered CCD im-

ages: 2006lv by G. Duszanowicz (Akersberga, Sweden, 0.32-m reflector); 2006mo and 2006mq by N. Joubert, N. Lee, and W. Li (LOSS/KAIT; cf. IAUC 8770); 2006mp by T. Puckett and R. Gagliano (0.60-m refl.; cf. IAUC 8762); and 2006mr and 2006ms by L. A. G. Monard (cf. IAUC 8731).

SN  2006 UT  α_{2000}  δ_{2000}  Mag.  Offset
2006lv Oct. 28.12 11^{h}32^{m}03.^{s}30 +36^{o}42′03′′.6 17.0 10′′ E, 12′′ N
2006mo Nov. 1.24 0 46 38.55 +36 19 57.8 18.2 2′′.4 W, 4′′.7 N
2006mp Nov. 3.99 17 12 00.20 +46 33 20.8 16.9 14′′.6 E, 2′′.7 S
2006mq Oct. 22.56 8 06 12.39 −27 33 45.4 12.7 16′′.6 E, 124′′ S
2006mr Nov. 5.76 3 22 42.84 −37 12 28.5 15.6 14′′ E, 0′′.3 N
2006ms Nov. 6.74 20 38 20.52 −52 06 55.9 17.3 2′′ E, 18′′ S

Additional R-band magnitudes for 2006lv in UGC 6517, by Duszanowicz unless noted otherwise: 2005 Apr. 5.86 UT, [18.0; 11.86, [18.0; 2006 Nov. 2.10, 16.7; 3.03, 15.8 (image by F. Andersson, 0.90-m reflector, near Uppsala; measurements by J. Warell). SN 2006lv appears to be a type-Ib/c supernova after maximum light (details on CBET 722). Additional KAIT magnitudes for 2006mo in MCG +06-02-17: Oct. 20.25, [19.0; Nov. 4.19, 18.2. MCG +06-02-17 and MCG +06-02-16 (UGC 480) are interacting galaxies. SN 2006mo is a type-Ia supernova discovered before maximum (details on CBET 722). Additional magnitudes by Puckett for 2006mp in MCG +08-31-29: Oct. 2, 14, and 24, [19.0; Nov. 5.01, 16.7. SN 2006mp is a type-Ia supernova before or at maximum light upon discovery (details on CBET 724). Additional KAIT magnitudes for 2006mq in ESO 494-G26: Feb. 24.12, [19.0; Nov. 5.50, 13.2. SN 2006mq is a type-Ia supernova well past maximum light (cf. CBET 731). Additional magnitudes for 2006mr in NGC 1316: Sept. 27.128, [17.0; Nov. 6.765, 15.2 ± 0.2. Note that SN 2006dd in NGC 1316 has position end figures 41^{s}.62, 13′′.0 (cf. IAUC 8723). SN 2006mr is a type-Ia supernova (details on CBET 729). Additional magnitudes for 2006ms in NGC 6935: (Digitized Sky Survey, red), [20.5; 2006 Oct. 12.81, [18.5; Nov. 7.735, 17.4 ± 0.2.

Also, 23 new supernovae (2006lw–2006mn via ESSENCE, with details on CBET 717; 2006mt–2006mx via SDSS II, from J. Frieman, with details on CBET 726) of mag 21 or fainter have been reported, all apparently of type Ia except for 2006lz (type II). SN 2006lu (cf. IAUC 8770) is a type-Ia supernova, a few days after maximum on Nov. 3 (details on CBET 716).

2006 November 10  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green